OpenStudy (anonymous):

Which of the following is the solution of log x – 5 0.001 = –3 ?

OpenStudy (anonymous):

the x-5 in small behind the log

OpenStudy (anonymous):

could you use the equation tool to clarify what the problem is?

OpenStudy (anonymous):

I don't know what the equation tool is?

OpenStudy (anonymous):

\[(x-5)\log(0.001) = -3\]is that your question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay, well when you have log, just written like that, it's in base 10. So first, calculate what's the log of 0.001 in base 10.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

okay, so then you have -3(x-5) = -3 so it's just a simple problem of finding x :)

OpenStudy (anonymous):

@sarahcoop3r did you get it?

OpenStudy (anonymous):

I got -4

OpenStudy (anonymous):

which isn't an option for me....

OpenStudy (anonymous):

oh. weird.. what are your options?

OpenStudy (anonymous):

5,7,13 and 15

OpenStudy (anonymous):

wait, how did you get 4? you just do: (x-5)=(-3/-3) so x=6. weird. i dunno, unless the question's been written wrong, my best guess would be to put in 5 as the answer.